OpenStudy (anonymous):

Which of the following is equivalent to |x + 11| = 13?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

What are your answer choices?

OpenStudy (anonymous):

x + 11 = 13 x + 11 = - 13 x + 11 = 13 or x + 11 = -13 - x + 11 = 13

OpenStudy (anonymous):

I would say it was your first choice.